I am trying to determine whether or not this statement makes sense and please explain.
"Gas prices jumped by 1% in January and 1% in February, so the increase in gas prices over the two months was 2%."

Found 2 solutions by ikleyn, richwmiller:
Answer by ikleyn(52864)

.

The statement is INCORRECT.


The increase in gas prices over the two months was


    ((1+0.01)*(1+0.01)-1)*100 = 2.01%.


Answer by richwmiller(17219)

in January the price rose from 100% of 10.0 to 101%
1.01*10=10.1
in February the price rose another 1% from 100% of 10.1 to 101% of 10.1
1.01*10.1=10.201
so the increase was 2.01 not a lot more than 2 but not 2 either